OpenStudy (anonymous):

Solve for x: 4x(x-5)-5x(x-4)=-1

OpenStudy (saifoo.khan):

x = 1, -1

OpenStudy (anonymous):

\[4x ^{2} -20x -5x^2 + 20x = -1 \] \[-x^2\] = -1 -x^2/-1 = -1/-1 x^2 = 1 \[{x} = \sqrt{1}\] REstrictions. X cannot = 5 or 4

OpenStudy (anonymous):

so\[x = \pm 1\]

OpenStudy (anonymous):

X= -1 and X= 1